OpenStudy (anonymous):

what are the dimensions of the lightest open-top right circular cylindrical can that will hold a volume of 1728 cm^3

OpenStudy (tkhunny):

"lightest"? I'm thinking you'll have to specify the materials and their weights.

OpenStudy (triciaal):

the question asked for the dimensions of the cylinder given the volume the material should not matter V = pi*r^2*h however the smallest height would need to be considered since we have 2 unknowns. except for r= 1 and h = 1 then r^2 will always be > h. r is the radius and h is the height it is odd to ask for the lightest dimension

OpenStudy (triciaal):

when the radius = height v = pi*r^3 r = (1728/pi)^1/3